The most recent financial statements for Bello Co. are shown here: Income Statement Sales Costs Taxable Balance Sheet 12,060 Debt $16,780 14,200 Fixed assets 34,200 Equity 29,480 $20,700 Current assets 6,500 Tot $46,260 Total $46,260 income Taxes (24%) 1,560 Net income $4,940 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 40 percent dividend payout ratio. What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, eg., 32.16.) Sustainable growth rate
Answer #1

The sustainable growth rate can be calculated as :

growth rate = ROE *b/ 1 -(ROE*b)

The return on equity = net income/ equity

=$4,940/ $29,480

= 16.76%

The dividend payout ratio is 40%, so, the retention ratio is (1 - dividend payout ratio) = 60%

So, the sustainable growth rate is = 0.6*0.1676/ 1 - (0.6*0.1676)

=0.1006/ 0.8994

=11.1852%

or, 11.18%.